OpenStudy (anonymous):

What is the square root by using imaginary number i for each of these. -7 -15 -81 -50 I got -81 but I'm not so sure if it's correct I got √ -81= 3i√ 9 but I'm not sure if this is correct of if I have to make the 9 a 3 as well.

OpenStudy (anonymous):

\[\sqrt{-81}=\sqrt{81\times (-1)}=\sqrt{81}\sqrt{-1}=9i\]

OpenStudy (anonymous):

the others are not perfect squares, so they will contain radicals for example \[\sqrt{-7}=\sqrt{7}i\]